Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
palma, taranto, bersanetti, gherardini - 16:20 Thursday 29 February 2024 (63445) Print this report
ETMs RHs powers step back

After the FSR scan performed by operator from 8:27UTC to 8:43UTC, we decided in agreement with the commissioning coordinator to completly revert the ETMs RHs step made on tuesday (63426) and restore the power settings of the ETMs RHs from Monday afternoon (63408) which currently represent the TCS configuration ensuring the operability of the ITF.

Thus, at 09.49 UTC we adjusted the power settings of the ETMs RHs as follows:

NE RH: 4.8 W → 4.7 W (13.2 V)

WE RH: 6.4 W → 6.3 W (15.3 V)

Unfortunately, we were unable to observe the effects of this adjustment due to issues encountered during the lock acquisition process caused by high wind activity at the site.

Search Help